OpenStudy (anonymous):

City Community College plans to increase its enrollment capacity to keep up with an increasing number of student applicants. The college currently has an enrollment capacity of 2200 students and plans to increase its capacity by 70 students each year. What will the enrollment capacity be 3 years from now? a. 2480 b. 2410 c. 2340 d. 2240

OpenStudy (anonymous):

2200,2200+70,2270+140 so the third year we have 2270+140=2410

OpenStudy (texaschic101):

2200 + 70 = 2270 (year 1) 2270 + 70 = 2340 (year 2) 2340 + 70 = 2410 (year 3) Jonask is correct :)

OpenStudy (anonymous):

Thanks ya'll